OpenStudy (error1603):

HELP! sin x + cos x = 1/2

OpenStudy (error1603):

Never mind. I got it. (sin x + cos x)^2 = Sin^2 x + Cos^2 x + 2 Cos x Sin x = 1 + 2 cos x sin x = 1/4 cos x sin x = -3/8 tan x + cot x = sin x/cosx + cos x/sin x = (sin^2 x + cos^2 x)/(sin x cos x) = 1/(sin x cos x) = 1/(-3/8) = -8/3 = -2.66666